Find s and r if (x+3) is a factor of sx^3 + rx^2 -28x +15 and has a remainder of -60 when divided by(x-3)


1
Expert's answer
2021-03-05T01:04:36-0500

Let us find "s" and "r" if "(x+3)" is a factor of "sx^3 + rx^2 -28x +15" and has a remainder of "-60" when divided by "(x-3)".


By the polynomial remainder theorem (little Bézout's theorem), the remainder of the division of a polynomial "\displaystyle f(x)=sx^3 + rx^2 -28x +15" by a linear polynomial "\displaystyle x-a" is equal to "\displaystyle f(a)". Therefore, we have the following system of linear equations


"\begin{cases}
s(-3)^3 + r(-3)^2 -28(-3) +15=0\\
s\cdot 3^3 + r\cdot 3^2 -28\cdot 3 +15=-60
\end{cases}"


which is equivalent to the following systems:


"\begin{cases}
-27s + 9r+99=0\\
27s + 9r-9=0
\end{cases}"


"\begin{cases}
-27s + 9r+99=0\\
18r+90=0
\end{cases}"


"\begin{cases}
-27s =-9r-99\\
r=-5
\end{cases}"


"\begin{cases}
-27s =-54\\
r=-5
\end{cases}"


"\begin{cases}
s =2\\
r=-5
\end{cases}"


Answer: "s=2,\ r =-5."